OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2 solve for g(f(3)). 5 7 9 10

zepdrix (zepdrix):

\[\Large\bf\sf f(\color{orangered}{x})\quad=\quad 2(\color{orangered}{x})+1\]\[\Large\bf\sf f(\color{orangered}{3})\quad=\quad 2(\color{orangered}{3})+1\] \[\Large\bf\sf g[\color{royalblue}{x}]\quad=\quad \frac{3[\color{royalblue}{x}]-1}{2}\] \[\Large\bf\sf g[\color{royalblue}{f(3)}]\quad=\quad \frac{3[\color{royalblue}{f(3)}]-1}{2}\]\[\Large\bf\sf g[\color{royalblue}{f(3)}]\quad=\quad \frac{3[\color{royalblue}{2(3)+1}]-1}{2}\]

OpenStudy (anonymous):

so 5?

zepdrix (zepdrix):

Hmm where is the 5 coming from? :o

OpenStudy (anonymous):

5 7 9 10

OpenStudy (anonymous):

the choices

zepdrix (zepdrix):

Why did you guess 5 though..? Was it just a random guess?

OpenStudy (anonymous):

no i did the math

OpenStudy (anonymous):

atleast i believe i did it correctly

zepdrix (zepdrix):

Hmm no that doesn't seem correct :o \[\Large\bf\sf g[\color{royalblue}{f(3)}]\quad=\quad \frac{3[\color{royalblue}{2(3)+1}]-1}{2}\]Simplify the blue stuff before anything else.\[\Large\bf\sf g[\color{royalblue}{f(3)}]\quad=\quad \frac{3[\color{royalblue}{7}]-1}{2}\]

OpenStudy (anonymous):

oh ok i see now

OpenStudy (anonymous):

it should be 10 then

OpenStudy (anonymous):

thank you

zepdrix (zepdrix):

good job \c:/